what is the molar mass of the solid acid HX, if 0.489 g of the acid is neutralized by 39.63 ml of 0.250 m naoh?
1
Expert's answer
2013-01-16T11:34:20-0500
The reaction between acid HX and NaOH:

NaOH + HX = NaX + H2O

n of NaOH = V*C/1000 = 39.63*0.250 / 1000 = 0.00991 mol

Mw = m/n = 0.489/0.00991 = 49